# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

### v2

@@ 4.1 Theoretical, managerial, and methodological implications
This study advances theories of teamwork, collaboration, and human-AI interaction by demonstrating how GPT-based AI agents reshape work processes, communication dynamics, and task execution in collaborative settings (Schneider et al., 2021). Our findings reveal that AI agents shift communication from social and emotional exchanges to task-oriented interactions, reducing interpersonal coordination costs and enabling participants to prioritize content generation. This aligns with theoretical frameworks suggesting that AI can streamline team processes by automating routine or procedural tasks (Wilson and Daugherty, 2018), but it extends these models by uncovering trade-offs in creative, multimodal workflows. Specifically, while AI agents enhanced text quality in ad creation, they underperformed in image-related tasks, suggesting that theories of human-AI collaboration must account for task-specific competencies and limitations, particularly in domains requiring diverse cognitive and creative skills (Dell’Acqua et al., 2023).
@@
Moreover, our work challenges the traditional conceptualization of AI as a passive tool or assistant (Noy and Zhang, 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024) by positioning GPT-based agents as active collaborators that shape the structure and quality of team outputs (Anthony et al., 2023, Makarius et al., 2020). For example, in our experiment, the AI’s proficiency in generating ad copy influenced participants to focus on generating text through instructions rather than direct edits. These changes in work processes suggest a need to reframe human-AI interaction theories to emphasize AI’s agency in collaborative processes. Rather than merely augmenting human efforts, AI agents co-create outcomes, introducing new dynamics that affect team performance.
Our findings also contribute to the literature on team cognition and distributed expertise (Collins et al., 2024). In human-human teams, expertise is often shared through social negotiation and emotional rapport, which fosters trust but can slow decision-making. In contrast, human-AI teams in our study bypassed much of this social overhead. This shift raises theoretical questions about how cognitive load and role allocation differ when humans collaborate with machines instead of other humans. Future theoretical work could explore these dynamics in other creative domains, such as product design, to develop a more nuanced understanding of how AI agents influence cognitive and collaborative processes across varied task environments.
@@
Our work also offers actionable insights for managers and organizations seeking to integrate GPT-based AI agents into collaborative workflows, particularly in creative and productivity-driven environments. Our findings demonstrate that AI agents significantly reduce social coordination costs, allowing team members to focus on task execution and content generation. Managers in industries like marketing, content creation, or consulting can leverage AI agents to streamline routine tasks, such as drafting campaign materials or reports, freeing human employees for strategic activities like audience analysis or brand positioning.
@@
However, the study also highlights critical limitations that managers must address to fully harness AI’s potential in multimodal workflows. While GPT-based agents excelled in text generation, they underperformed in image-related tasks. This trade-off supports the idea that organizations should adopt a hybrid approach along a jagged edge of AI capabilities (Dell’Acqua et al., 2023), allocating tasks within AI capabilities to agents and those outside AI capabilities, such as image selection, to humans. Such an approach can mitigate the quality disparities observed in our study and are particularly relevant in domains where compelling visual elements drive downstream outcomes.
@@
Furthermore, our results provide a framework for structuring human-AI teams to optimize productivity and performance. AI’s ability to handle bulk editing tasks—as evidenced by the 71% fewer direct copy edits in human-AI teams—allowed participants to engage in more content ideation and quality control. Managers can apply this division of labor to design efficient workflows, assigning AI agents to repetitive or data-intensive tasks while reserving human expertise for creative oversight and final approvals. A practical example might involve a content production team where AI drafts blog posts or social media captions, which editors then polish for tone and brand alignment. This approach not only boosts efficiency but also leverages human strengths in contextual judgment, which AI currently lacks, particularly in nuanced visual tasks. To implement such workflows, organizations should invest in training programs that teach employees how to delegate tasks to AI effectively, fostering a collaborative culture that maximizes both technological and human contributions.
Finally, the study underscores the importance of aligning AI deployment with organizational goals and team dynamics. The comparable team-level productivity between human-AI and human-human teams, despite human-AI teams having half the human resources, suggests that AI can serve as a near-substitute for additional personnel in certain contexts. Managers could use this insight to scale team output without proportionally increasing headcount, a cost-effective strategy for startups or resource-constrained firms. However, they must remain mindful of the social dynamics altered by AI integration. The decrease in social and emotional messages in human-AI teams indicates less time spent on rapport-building, which could impact team cohesion in long-term projects. Managers should monitor team morale and consider hybrid models—such as combining AI assistance with periodic human-human collaboration—to maintain a balance between efficiency and interpersonal connection.
@@
This study advances the methodological landscape for studying human-AI collaboration through the development of the Pairit platform, a novel experimental framework that redefines how researchers investigate team dynamics in real-time, task-driven settings. Unlike traditional approaches that rely on aggregate performance metrics or isolated AI interactions (Noy and Zhang, 2023, Dell’Acqua et al., 2023), Pairit captures granular, time-stamped data on every keystroke, message, edit, and API call, enabling a detailed reconstruction of collaborative workflows. In our experiment, this capability allowed us to analyze 183,691 messages, 1,960,095 text edits, and 63,656 image edits, revealing nuanced patterns such as how AI-driven suggestions shifted human effort from direct text editing to instructional messaging. This methodological advance allows us to analyze both aggregate and detailed dynamics of human-AI teamwork and offers researchers new insights into decision-making, task allocation, and interaction processes.
@@
Pairit’s multimodal capabilities further distinguish it as a methodological tool, supporting tasks that integrate text, images, and external API interactions. While our study focused on ad creation, Pairit’s architecture is extensible to diverse domains, such as software development (e.g., AI generating code while humans debug), scientific writing (e.g., AI drafting literature reviews), or multimedia content production (e.g., designing virtual reality environments). This versatility allows researchers to explore human-AI collaboration in tasks requiring varied cognitive and creative skills, addressing a gap in prior studies that often limit AI to text-based functions (Brynjolfsson et al., 2023). For example, a future study could use Pairit to analyze how AI agents assist in architectural planning, tracking both textual specifications and visual renderings to assess collaborative efficiency and output quality. By capturing multimodal interactions, Pairit provides a comprehensive framework for studying the interplay of different task components, a methodological leap forward for understanding complex workflows.
These methodological contributions have broader implications for experimental research and organizational practice. Pairit’s ability to simulate realistic collaborative environments offers a testing ground for refining theoretical models of human-AI interaction and informing the design of AI-driven workflows. For instance, researchers could leverage the platform to study how AI agents perform in agile software development sprints, using time-stamped data to optimize task assignments and evaluate productivity trade-offs. Ultimately, Pairit’s methodological innovations bridge the gap between controlled experiments and real-world applications and provides a robust, scalable framework for advancing our understanding of human-AI collaboration.
@@ 4.2 Limitations and future directions
While we conducted one of the first truly agent-randomized experiments in human-AI collaboration, combining laboratory and field tests, on a newly developed platform capable of recording fine-grained collaboration dynamics, our work is not without limitations. First, while we observe substantial differences in communication patterns between human-human and human-AI teams, the underlying mechanisms or explanations for these differences are not always obvious. They could stem from the nature of AI communication, which tends to prioritize content-related messages, or from human participants adjusting their behavior upon realizing their partner is an AI. Our exit surveys found that human participants were good at predicting whether their partner was an AI (2.7 points higher on a 7-point Likert scale, $p<.001$), suggesting AI communication and human adaptation are both at work in these mechanisms. But, future work is needed to elucidate these potential explanations and better understand the drivers of these shifts. Second, the limitations of current AI models, particularly Vision GPT models, present challenges for multimodal tasks. These models are optimized for next-word generation and specific visual tasks, such as identifying items in images, but are not designed for nuanced assessments like image quality prediction. This limitation likely contributed to the lower image quality observed in human-AI teams and underscores the need for AI systems tailored to specific creative and evaluative tasks. However, AI is a moving target and image selection and design competency likely will improve over time. As with any study of AI, results may drift as AI changes. Moreover, while our study provides a controlled experimental context, it may not fully capture the complexities of long-term collaboration with AI agents in real-world environments. Future research should explore how these dynamics evolve over extended periods and in more diverse task domains to validate and expand on our findings.
@@
Our work opens several promising avenues for future research. First, the limitations of current Vision GPT models suggest a need for separate, specialized image generation models. Future studies could incorporate state-of-the-art visual models designed specifically for tasks like image quality prediction and generation, enabling a more nuanced understanding of AI contributions to multimodal creative tasks. Second, our findings are grounded in the context of ad design, but they invite exploration across a range of other collaborative domains. Important contexts include software development (e.g., coding), data analysis, collaborative writing, and financial accounting. Investigating Human-AI collaboration in these areas could reveal domain-specific dynamics and inform the broader design of AI systems tailored to various professional workflows. Finally, extending this research to longitudinal settings could provide insights into how Human-AI collaboration evolves over time. Long-term studies could examine the sustainability of productivity gains, the development of trust, and the potential for “learning effects” where humans adapt to working with AI agents or vice versa. These investigations would bridge the gap between controlled experimental settings and real-world applications, enhancing the generalizability of our findings. We intend to make Pairit freely available to academics advancing peer-reviewed research into human-AI collaboration. We believe this platform could support hundreds, if not thousands, of new experimental studies on the effects of AI on productivity, performance, and work processes.
@@ 4.3 Conclusion
This study provides novel insights into how AI agents reshape work processes, productivity, and performance in collaborative settings. Our findings reveal that collaborating with AI agents significantly increases communication while decreasing (but not completely eliminating) direct engagement with the work product. Human-AI teams focus more on content and process-oriented communication, reducing the need for social and emotional exchanges compared to human-human teams. Humans in human-AI teams experienced 73% greater productivity per worker, suggesting that AI agents can act as near-substitutes for additional human collaborators. However, AI had countervailing effects on different dimensions of output quality—while ad text quality improved in human-AI teams, image quality declined, highlighting a trade-off in multimodal tasks. Field evaluations of ad effectiveness showed that ads with higher image quality (produced by human-human collaborations) and higher text quality (produced by human-AI collaborations) performed significantly better on click-through rates, view through rates, and cost per click metrics.
A key innovation of this study is the introduction of Pairit, an experimental platform that enables real-time collaboration between humans and multimodal AI agents. By providing researchers with a controlled environment to study these dynamics, Pairit opens new avenues for understanding how AI agents can enhance teamwork and productivity. Its versatility allows for the exploration of a wide range of collaborative tasks beyond ad creation, including software development, data analysis, and collaborative writing, making it a valuable tool for researchers across disciplines. As AI agents become increasingly integrated into modern workflows, platforms like Pairit will be essential for advancing our knowledge of human-AI collaboration and designing more effective human-AI partnerships.
@@ 5 Acknowledgments
A Prompts for message labeling

### v3

@@ 5.1 Theoretical, managerial, and methodological implications
This study extends teamwork theory to human-AI collaboration, demonstrating how LLM-based AI agents reshape work processes, communication dynamics, and task execution in collaborative settings (Schneider et al., 2021). Our findings connect to three effects documented in the AI literature: productivity gains (Noy and Zhang, 2023, Brynjolfsson et al., 2023), the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), and diversity collapse (Padmakumar and He, 2024, Chen and Chan, 2024). We find that human-AI collaboration induces a delegation workflow—task-oriented, directive communication and increased delegation to AI—which improves text quality (an AI strength) but not image quality (an AI weakness), with treatment effects attenuated by these jagged frontier effects. Delegation also reduces output diversity, as teams that delegate more produce outputs more similar to the typical human-AI output. This extends theoretical frameworks on human-AI collaboration design (Agarwal et al., 2025) and behavioral trust in AI teammates (Zhang et al., 2023) from decision-making contexts to agentic, multimodal creative work.
The shift from interpersonal to task-oriented communication aligns with work showing that humans apply social behaviors to machines (Reeves and Nass, 1996), but that these behaviors differ systematically from human-human interactions in that people interacting with AI become more demanding and instrumental (Tey et al., 2024). This shift reflects an optimal adaptation to AI partners that follow instructions reliably and require no social maintenance. While AI can streamline team processes (Wilson and Daugherty, 2018), our findings revealed trade-offs. In our experiment, use of AI enhanced ad text quality but reduced image quality. Our field experiment also revealed that these quality differences map to distinct performance outcomes. Higher text quality was associated with higher click-through rates and view-through duration, while higher image quality was associated with reductions in cost-per-click. This suggests the jagged frontier extends beyond output quality to ad performance in the field as AI capabilities create comparative advantages in different outcome metrics. Theories of human-AI collaboration must therefore account not only for task-specific competencies but also for how these competencies translate to heterogeneous downstream outcomes (Dell’Acqua et al., 2023).
Critically, we identify recognition, or participants’ awareness that their partner is an AI, as a key moderator of these dynamics. Participants who correctly identified their AI partner shifted toward more task-oriented communication and delegated more work, suggesting that awareness of AI capabilities triggers behavioral adaptation. This demonstrates that recognition is not merely perceptual but consequential: it reshapes collaboration behavior in ways that affect output quality. The asymmetry in recognition accuracy (94.5% in human-AI teams versus 57.6% in human-human teams) suggests that AI communication patterns are distinctive and detectable, which may itself shape how humans calibrate their collaborative strategies.
@@
Moreover, our work challenges the traditional conceptualization of AI as a passive tool or assistant (Noy and Zhang, 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024) by positioning LLM-based agents as active collaborators that shape team dynamics and, in turn, productivity and output quality (Anthony et al., 2023, Makarius et al., 2020). For example, the AI’s proficiency in generating ad copy influenced participants to focus on instructions rather than direct edits. Rather than merely augmenting human efforts, AI agents co-create outcomes, introducing new dynamics that affect team performance.
Our findings also contribute to the literature on team cognition and distributed expertise (Collins et al., 2024). In human-human teams, expertise is often shared through interpersonal exchanges, which foster trust but can slow decision-making. In contrast, human-AI teams in our study bypassed much of this social overhead. Interestingly, we found that interpersonal communication predicts higher output diversity, suggesting that social exchanges may facilitate the sharing of novel ideas between collaborators. This creates a tension: the efficiency gains from task-oriented communication come at the cost of creative variation. We show that traditional teamwork dimensions operate differently in human-AI teams: coordination becomes directive rather than mutual, conflict is avoided rather than resolved, and task allocation becomes delegation rather than negotiation. These patterns suggest that teamwork theory remains applicable to human-AI collaboration but requires adaptation for the asymmetry inherent in these teams.
@@
Our work also offers actionable insights for managers and organizations seeking to integrate AI agents into collaborative workflows, particularly in creative and productivity-driven environments. Our findings demonstrate that AI agents significantly reduce social coordination costs, allowing team members to focus on task execution and content generation. Managers in industries like marketing, content creation, or consulting can leverage AI agents to streamline routine tasks, such as drafting campaign materials or reports, freeing human employees for strategic activities like audience analysis or brand positioning.
@@
However, the study also highlights critical limitations that managers must address to fully harness AI’s potential in multimodal workflows. While LLM-based agents excelled in text generation, they underperformed in image-related tasks. This trade-off supports the idea that organizations should adopt a hybrid approach along the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), allocating tasks within AI capabilities to agents and those outside AI capabilities, such as image selection, to humans or fine-tuning models to perform better on those tasks. Such an approach can mitigate the quality disparities observed in our study and is particularly relevant in domains where compelling visual elements drive downstream outcomes.
@@
Furthermore, our results provide a framework for structuring human-AI teams to optimize productivity and performance. AI’s ability to handle bulk editing tasks allowed participants to engage in more content ideation and quality control. Managers can apply this division of labor to design efficient workflows, assigning AI agents to repetitive or data-intensive tasks while reserving human expertise for creative oversight and final approvals. To implement such workflows, organizations should invest in training programs that teach employees how to delegate tasks to AI effectively, fostering a collaborative culture that maximizes both technological and human contributions.
Finally, the study underscores the importance of aligning AI deployment with organizational goals and team dynamics. Our findings reveal that higher delegation to AI reduces output diversity, a concern for organizations that value creative variation. To mitigate diversity collapse, managers could design workflows that preserve human-driven ideation, such as human-only brainstorming sessions before AI-assisted execution. Although human-AI teams produce fewer total submissions than human-human teams, the 50% increase in individual productivity suggests that AI can serve as a partial substitute for additional personnel in certain contexts. Managers could use this insight to scale team output without proportionally increasing headcount, a cost-effective strategy for startups or resource-constrained firms. However, they must also be mindful of “collapsing onto the jagged frontier”: as teams rely more on AI, they may improve at tasks where AI excels (text generation) while atrophying at tasks where AI underperforms (image selection). Organizations should ensure that human skills in domains in which AI is weak are maintained through deliberate practice or task rotation.
@@
This study advances the methodological landscape for studying human-AI collaboration through the development of the Pairit platform, a novel experimental framework that redefines how researchers investigate team dynamics in real-time, task-driven settings. Unlike traditional approaches that rely on aggregate performance metrics or isolated AI interactions (Noy and Zhang, 2023, Dell’Acqua et al., 2023), Pairit captures granular, time-stamped data on every keystroke, message, edit, and API call, enabling a detailed reconstruction of collaborative workflows. In our experiment, this capability allowed us to analyze 182,607 messages, 1,889,559 text edits, 62,119 image edits, and 10,074 AI-generated images, revealing nuanced patterns such as how AI-driven suggestions shifted human effort from direct text editing to instructional messaging. This methodological advance allowed us to analyze both aggregate and granular dynamics of human-AI teamwork and offers researchers new opportunities to discover insights regarding human-AI decision-making, task allocation, and interaction processes.
@@
Pairit’s multimodal capabilities further distinguish it as a methodological tool, supporting tasks that integrate text, images, and external API interactions. While our study focused on ad creation, Pairit’s architecture is extensible to diverse domains, such as software development (e.g., AI generating code while humans debug), scientific writing (e.g., AI drafting literature reviews), or multimedia content production (e.g., designing virtual reality environments). This versatility allows researchers to explore human-AI collaboration in tasks requiring varied cognitive and creative skills, addressing a gap in prior studies that often limit AI to text-based functions (Brynjolfsson et al., 2023).
These methodological contributions have broader implications for experimental research and organizational practice. Pairit’s ability to simulate realistic collaborative environments offers a testing ground for refining theoretical models of human-AI interaction and informing the design of AI-driven workflows. Ultimately, Pairit’s methodological innovations bridge the gap between controlled experiments and real-world applications and provide a robust, scalable framework for advancing our understanding of human-AI collaboration.
@@ 5.2 Limitations and future directions
While we conducted one of the first truly agent-randomized experiments in human-AI collaboration, combining laboratory and field tests, on a newly developed platform capable of recording fine-grained collaboration dynamics, our work is not without limitations. First, although we identify task-orientation, delegation, and recognition as mechanisms underlying human-AI collaboration dynamics, we cannot fully disentangle whether these patterns stem from the nature of AI communication or from human participants adapting their behavior upon recognizing their partner as AI. Our mediation analyses tracing the pathway from recognition to communication to delegation to quality are post hoc and correlational; while they are consistent with a causal pathway, alternative explanations remain possible. For instance, participants who are more comfortable with AI may both recognize it more readily and adopt more efficient collaboration strategies, with no causal relationship between recognition and behavior. Second, the limitations of current AI models, particularly vision models, present challenges for multimodal tasks. These models are optimized for next-word generation and specific visual tasks, such as identifying items in images, but are not designed for nuanced assessments like image quality prediction. This limitation likely contributed to the lower image quality observed in human-AI teams and underscores the need for AI systems tailored to specific creative and evaluative tasks. However, AI is a moving target and image selection and design competency likely will improve over time. As with any study of AI, results may drift as AI changes. Moreover, while our study provides a controlled experimental context, it may not fully capture the complexities of long-term collaboration with AI agents in real-world environments. Future research should explore how these dynamics evolve over extended periods and in more diverse task domains to validate and expand on our findings.
@@
Regarding generalizability, our findings should apply to tasks that share key features with ad creation: multi-modal creative tasks requiring both text and visual elements, where AI capabilities vary across modalities (the jagged frontier); iterative production tasks involving multiple rounds of generation, evaluation, and refinement; and settings with objective quality metrics. Our findings may be less applicable to high-stakes decision tasks where errors are costly and irreversible or tasks requiring deep domain expertise where AI capabilities are more limited. While our study captured within-session adaptation (humans learning AI capabilities through multiple ad iterations), longitudinal dynamics across repeated collaborations remain unexplored. The core mechanism we identify (task-oriented communication, higher delegation, and capability-dependent quality effects) should generalize broadly, but the direction of quality effects will depend on the specific task and AI capabilities.
@@
Our work opens several promising avenues for future research. First, ours is one of the first randomized controlled trials to evaluate how AI agents affect work and communication processes. In our detailed, task-level analysis, we uncovered specific effects on communication, collaboration, and work processes, but these findings represent only a first step. Future work should provide more task-level evidence about how collaboration with AI agents changes communication, teamwork, work processes, and organization design, and how these effects drive productivity and performance across task and organizational contexts. Such research is essential to understanding the impact of AI on work and organizational performance. Second, the limitations of current vision models suggest a need for separate, specialized image generation models. Future studies could incorporate state-of-the-art visual models designed specifically for tasks like image quality prediction and generation, enabling a more nuanced understanding of AI contributions to multimodal creative tasks. Third, our findings are grounded in the context of ad design, but they invite exploration across a range of other collaborative domains. Important contexts include software development (e.g., coding), data analysis, collaborative writing, and financial accounting. Investigating Human-AI collaboration in these areas could reveal domain-specific dynamics and inform the broader design of AI systems tailored to various professional workflows. Finally, extending this research to longitudinal settings could provide insights into how Human-AI collaboration evolves over time. Long-term studies could examine the sustainability of productivity gains, the development of trust, and the potential for “learning effects” where humans adapt to working with AI agents or vice versa. These investigations would bridge the gap between controlled experimental settings and real-world applications, enhancing the generalizability of our findings. We intend to make Pairit freely available to academics advancing peer-reviewed research into human-AI collaboration. We believe this platform could support hundreds, if not thousands, of new experimental studies on the effects of AI on productivity, performance, and work processes.
@@ 5.3 Conclusion
Our work provides novel insights into how AI agents reshape work processes, productivity, and performance in collaborative settings. Our findings connect to three effects documented in the AI literature. First, human-AI teams achieved 50% higher productivity per worker, suggesting that AI agents can substitute for human collaborators. Second, we observed the jagged frontier of AI capabilities: text quality improved in human-AI teams while image quality declined, with these offsetting effects explaining similar field performance. Third, we find diversity collapse: teams that delegated more to AI produced outputs more similar to the typical human-AI output. We then found three mechanisms underlying these effects: task orientation (human-AI teams sent 25% more task-oriented messages and 18% fewer interpersonal messages), delegation (participants delegated 17% more work to AI and made 62% fewer direct edits), and recognition (participants who correctly identified their AI partner were more task-oriented and delegated more). Together, these mechanisms constitute a delegation workflow in which humans direct and delegate work to AI collaborators to unlock performance gains. These findings suggest that the productivity, quality, and output diversity effects of AI depend critically on how humans communicate with, collaborate with, and delegate work to AI agents.
@@ 6 Acknowledgments
Appendix A Big Five Personality Balance Checks
To complement the demographic balance checks reported in Table 1, we examined whether Big Five personality traits were balanced across experimental conditions. Table A shows no significant differences in openness, conscientiousness, extraversion, agreeableness, or neuroticism between human-human and human-AI conditions, confirming successful randomization.
Trait All Human-AI Human-Human t-statistic (SE) p-value
Openness 0.71 $\pm$ 0.20 0.71 $\pm$ 0.21 0.72 $\pm$ 0.19 -0.869 (0.009) 0.385
Conscientiousness 0.79 $\pm$ 0.18 0.78 $\pm$ 0.18 0.79 $\pm$ 0.17 -1.039 (0.008) 0.299
Extraversion 0.56 $\pm$ 0.21 0.55 $\pm$ 0.21 0.56 $\pm$ 0.21 -0.949 (0.009) 0.343
Agreeableness 0.70 $\pm$ 0.22 0.70 $\pm$ 0.21 0.69 $\pm$ 0.23 0.473 (0.010) 0.636
Neuroticism 0.48 $\pm$ 0.23 0.49 $\pm$ 0.24 0.47 $\pm$ 0.22 1.589 (0.010) 0.112
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Personality traits are normalized from a 7-point Likert scale.
Table A: Big Five personality trait balance checks
Appendix B Productivity
Table B reports the full regression results for productivity outcomes at both team and individual levels. At the team level, human-AI teams produced fewer total submissions, which is expected given that these teams have half as many human workers. At the individual level, however, participants in human-AI teams submitted 50% more ads than their counterparts in human-human teams.
Team Individual
Intercept 7.623∗∗∗ 3.850∗∗∗ 4.601∗∗∗
(0.277) (0.129) (0.346)
Human-AI -1.847∗∗∗ 1.925∗∗∗ 1.954∗∗∗
(0.315) (0.197) (0.197)
Demographics No No Yes
Observations 1751 2234 2234
Table B: Productivity of human-AI and human-human teams.
Appendix C Work Actions
Table C presents detailed results on work actions across experimental conditions. Participants in human-AI teams made substantially fewer direct copy edits but engaged in more image edits and AI image generations, suggesting a shift from direct text manipulation to delegation and visual content creation.
Copy Edits Image Edits AI Images
Intercept 1676.458∗∗∗ 28.397∗∗∗ 6.067∗∗∗
(57.646) (1.559) (0.390)
Human-AI -1032.406∗∗∗ 6.592∗∗∗ 0.328
(33.588) (0.931) (0.219)
Demographics Yes Yes Yes
Observations 2234 2234 2234
Table C: Collaborating with AI reduces copy edits but increases image edits and AI image generations at the individual level.
Appendix D Message Categories
Table D reports the distribution of message categories across experimental conditions. Messages were classified into content, process, social, emotional, feedback, and other categories using GPT-4o-mini. Human-AI teams sent significantly more content- and process-oriented messages, while human-human teams sent more social and emotional messages.
Content Process Emotional Social Feedback
Intercept 0.179∗∗∗ 0.187∗∗∗ 0.195∗∗∗ 0.391∗∗∗ 0.048∗∗∗
(0.012) (0.010) (0.010) (0.013) (0.005)
Human-AI 0.048∗∗∗ 0.040∗∗∗ -0.044∗∗∗ -0.067∗∗∗ 0.022∗∗∗
(0.006) (0.006) (0.005) (0.007) (0.003)
Demographics Yes Yes Yes Yes Yes
Observations 2228 2228 2228 2228 2228
Table D: Collaborating with AI increases content-, process-, and feedback-related messages while decreasing social and emotional messages.
Appendix E Ad Copy Completion
Table E examines whether productivity differences might be driven by submission of incomplete ads. Participants in human-AI teams had consistently higher completion rates for all ad copy elements (headline, primary text, and description), indicating that the productivity gains reflect genuine output rather than incomplete submissions.
Headline Primary Text Description
Intercept 0.808∗∗∗ 0.769∗∗∗ 0.719∗∗∗
(0.025) (0.026) (0.027)
Human-AI 0.150∗∗∗ 0.169∗∗∗ 0.161∗∗∗
(0.014) (0.014) (0.015)
Demographics Yes Yes Yes
Observations 2234 2234 2234
Table E: Individuals in Human-AI teams submit ads with more copy completed.
Appendix F Survey Interface
Figure F shows the user interface for the ad quality survey described in Section 3.8.
Figure F: The user interface for the ad quality survey.
Appendix G Human Ratings with Full Covariates
Table G presents the complete regression results for human quality ratings, including demographic controls and work process covariates.
Text Image Click
Intercept 5.011∗∗∗ 5.045∗∗∗ 4.666∗∗∗ 4.413∗∗∗ 3.698∗∗∗ 3.539∗∗∗
(0.038) (0.048) (0.034) (0.045) (0.035) (0.044)
Human-AI 0.327∗∗∗ 0.432∗∗∗ -0.133∗∗∗ -0.165∗∗∗ -0.013 0.027
(0.024) (0.052) (0.021) (0.044) (0.021) (0.044)
# Submissions -0.028∗∗∗ -0.012∗∗∗ -0.018∗∗∗
(0.002) (0.001) (0.001)
# AI Images Generated 0.008∗∗∗ 0.029∗∗∗ 0.023∗∗∗
(0.002) (0.002) (0.002)
# Human Messages 0.004∗∗∗ 0.002∗∗∗ 0.002∗∗∗
(0.000) (0.000) (0.000)
# Human Copy Edits 0.140∗∗∗ 0.042∗∗ 0.080∗∗∗
(0.017) (0.014) (0.015)
# Human Copy Edits $\times$ (H-AI) -0.032 -0.001 -0.019
(0.027) (0.027) (0.026)
# Human Image Edits -2.165 -1.054 -0.664
(0.893) (0.704) (0.722)
# Human Image Edits $\times$ (H-AI) 2.255 2.072 1.832
(1.004) (0.822) (0.827)
Demographics Yes Yes Yes Yes Yes Yes
Observations 45803 45803 45803 45803 45803 45803
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Standard errors are clustered at the ad level. Copy and image edits are per thousand.
Table G: Human ratings of ads with full covariates.
Appendix H AI Ratings with Full Covariates
AI ratings.
To obtain AI ratings of ad quality, we prompt OpenAI’s multimodal gpt-4o-mini-2024-07-18, which supports image input and structured outputs. [^7] See OpenAI’s documentation on vision and structured outputs. To make AI evaluations comparable to human evaluations, we asked the same questions of the AI as those given to human evaluators (see Section 3.8). Each item was evaluated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Appendix L for the full prompt.
Results.
Table H presents the complete regression results for AI quality ratings, using the same evaluation criteria as human raters. The treatment effects on text and image quality remain robust after controlling for participant characteristics and the number of edits made during the collaboration.
Text Image Click
Intercept 5.463∗∗∗ 5.396∗∗∗ 6.494∗∗∗ 6.369∗∗∗ 5.281∗∗∗ 5.253∗∗∗
(0.026) (0.038) (0.026) (0.034) (0.022) (0.032)
Human-AI 0.128∗∗∗ 0.120∗∗∗ -0.011 -0.114∗∗∗ 0.071∗∗∗ 0.040
(0.015) (0.031) (0.015) (0.030) (0.013) (0.027)
# Submissions -0.013∗∗∗ -0.010∗∗∗ -0.008∗∗∗
(0.001) (0.001) (0.001)
# AI Images Generated 0.024∗∗∗ 0.029∗∗∗ 0.020∗∗∗
(0.001) (0.001) (0.001)
# Human Messages 0.003∗∗∗ 0.002∗∗∗ 0.001∗∗∗
(0.000) (0.000) (0.000)
# Human Copy Edits -0.004 -0.005 -0.025
(0.011) (0.010) (0.010)
# Human Copy Edits $\times$ (H-AI) -0.038 0.004 -0.061∗∗∗
(0.018) (0.018) (0.016)
# Human Image Edits -1.460∗∗ -1.909∗∗∗ -1.259∗∗
(0.527) (0.446) (0.425)
# Human Image Edits $\times$ (H-AI) 0.342 2.381∗∗∗ 0.999
(0.581) (0.507) (0.477)
Demographics Yes Yes Yes Yes Yes Yes
Observations 11024 11024 11024 11024 11024 11024
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Standard errors are clustered at the ad level. Copy and image edits are per thousand.
Table H: AI ratings of ads with full covariates.
Appendix I Field Experiment View Metrics
Table I reports the full regression results for view-through rate (VTR) and view-through duration (VTD) from the field experiment on X. While collaboration type showed no direct effect on view metrics, text quality was significantly associated with longer viewing duration, suggesting that the higher text quality produced by human-AI teams translates to greater audience engagement.
VTR VTD (log-sec)
Intercept 0.000 -0.000 -0.000 0.518∗∗∗ 0.000 0.000
(0.006) (0.012) (0.012) (0.053) (0.119) (0.121)
Human-AI 0.001 -0.000 0.012 0.002
(0.004) (0.004) (0.037) (0.038)
Click -0.000 -0.000 -0.008 -0.008
(0.003) (0.003) (0.027) (0.027)
Image -0.003 -0.003 -0.002 -0.002
(0.002) (0.002) (0.023) (0.023)
Text 0.002 0.002 0.037 0.037
(0.002) (0.002) (0.019) (0.019)
Spend 0.001∗∗∗ 0.001∗∗∗ 0.001∗∗∗ -0.001 0.004 0.004
(0.000) (0.000) (0.000) (0.002) (0.002) (0.002)
Campaign RE Yes Yes Yes Yes Yes Yes
Observations 4997 4997 4997 4997 4997 4997
  • Notes: p<<0.1, p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects.
Table I: Effects on view-through rate (VTR) and view-through duration (VTD) from the field study.
Appendix J Prompts for Message Labeling